# Global magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regime

### v1

@@ 4.2.1 Emergence and saturation of the MRI
Figure 4: Temporal evolution of the volume-averaged total $\alpha$ (black) and Maxwell $\alpha$ (red) in the MRI active zone, which are larger in the higher resolution simulation ZNF-HRES (dotted) compared to ZNF-FID (solid). The volume average is computed over the spherical wedge: $r\in[2,4]$, $\theta\in[z=-4H,z=4H]$ and $\phi\in[0,2\pi)$.
We begin by characterising the saturated MRI dynamo in the ideal MHD region of the disc. As shown in Fig. 4, the volume-averaged $\alpha$ in the ideal MHD zone exhibits rapid exponential growth, reaching saturation within approximately $3t_{\mathrm{DZI}}$, and remains sustained for the duration of simulation. The time-averaged value is $\alpha\sim 8\times 10^{-3}$, with the Maxwell component contributing $\alpha_{\mathcal{M}}\sim 5\times 10^{-3}$. This behaviour confirms the presence of a persistent MRI-driven dynamo operating in a fully developed turbulent state. For comparison, the global ZNF dead–active zone interface simulations with radiative transfer by Flock et al. (2017), report a higher volume-averaged stress of $\alpha\sim 3\times 10^{-2}$ in the MRI active zone of their fiducial model. However, their azimuthally restricted domain employed a higher effective resolution, equivalent to 2048 cells in azimuth.
To assess the sensitivity of this evolution to the numerical grid, we conduct a resolution test using ZNF-AZ. The disc structure, evolution, and large-scale magnetic features resemble those in ZNF-FID, but differences arise due to changes in the saturated strength of the ZNF MRI. As shown in Fig. 4, the total and Maxwell components of the saturated, volume-averaged $\alpha$, in the active zone, are about $60$ per cent higher in ZNF-HRES (dotted lines), enhancing angular momentum transport and accelerating viscous evolution by a similar factor, under thin-disc assumptions. This underscores the difficulty of extracting precise accretion rates from global ZNF MRI simulations at these resolutions. Whether higher-resolution global simulations reproduce the non-converging behaviour in stratified isothermal shearing boxes (see table 1 in Ryan et al., 2017), where $\alpha$ falls with increasing resolution, remains unclear.
@@ 4.2.2 Large-scale coherent poloidal magnetic field loops
In the MRI active region, large-scale poloidal field loops are generated self-consistently from an initial toroidal field (12) and grow to fill the extent of the active disc region. In Fig. 5, we present the magnetic field configuration in the inner disc, averaged in azimuth and over the time interval $t_{\mathrm{in}}\in[200,700]$, to enhance the visibility of the coherent field structures against the background turbulent state. The poloidal field configuration, $\langle\mathbf{B_{p}}\rangle_{\phi,t}$, is visualised using field lines (black) and a LIC overlay (greyscale). A set of concentric clockwise large-scale poloidal field loops is clearly visible. Moreover, once these structures form they persist throughout the extent of the simulation, as discussed in Section 6.
Previous work by Jacquemin-Ide et al. (2024) demonstrated the self-consistent generation of large-scale poloidal magnetic field loops in global MRI active simulations of accreting black holes. In contrast to our initial conditions (Section 3.3.1), their study began with a strong, uniform axisymmetric toroidal magnetic field with a plasma-beta of $\beta=5$ and a torus scale height of $H/r\sim 0.2$ in the bulk disc, corresponding to a relatively thick torus structure. However, they raised concerns regarding the efficiency of dynamo-driven field generation in thinner discs, which are less effective at advecting magnetic flux (Lubow et al., 1994). Therefore, this work provides evidence for the formation and persistence of similar large-scale poloidal field loops in a thinner disc with $H/R=0.1$, a regime more representative of protoplanetary discs. This lends tentative support to the notion that thin MRI active discs can generate and sustain large-scale magnetic fields self-consistently, without requiring a strong initial vertical-net flux.
Figure 5: Magnetic field configuration for ZNF-FID around the dead–active zone interface, averaged in azimuth and over the interval $t_{\text{in}}\in[200,700]$ to enhance the visibility of coherent field structures. The poloidal magnetic field $\langle\mathbf{B_{p}}\rangle_{\phi,t}$, shown using black field lines and a LIC overlay in greyscale, forms coherent clockwise loops spanning the extent of the MRI active region. The sign of the (normalised) toroidal field $R\langle B_{\phi}\rangle_{\phi,t}$ (background colour) near the disc surface layers (dashed white lines), is negatively correlated with the sign of the radial magnetic field.
@@ 4.3 Vertical structure of the active zone
4.3.1 Surface layer accretion
@@
In the MRI active, ideal MHD region of the disc ($R<1$ au) there is the rapid and sustained development of a vertical layered accretion structure, which we believe has not been seen before in the ZNF configuration. The majority of the accretion is localised just below the disc surface layers at $z\approx\pm 3H$, in a near symmetric manner across the disc midplane. To quantify the connection between accretion and magnetic stresses, we perform an extensive temporal average over the interval $t_{\mathrm{in}}\in[200,700]$ in this section.
@@
Figure 6: Spacetime $(\theta,t)$ diagrams for ZNF-FID, along the contour $r=0.75$ au for the azimuthally averaged toroidal field $\langle B_{\phi}\rangle_{\phi}$ (top) and Maxwell $\alpha$ component (22) (bottom). The disc–corona transition is at $z=\pm 4H$ (dashed black lines) and the surface layers above the MRI active disc have a tightly wound strong magnetic field configuration which drives strong local radial angular momentum transport.
@@
Figure 7: Vertical structure of the accretion flow in the MRI active region, quantified by $\langle\rho u_{r}\rangle_{\phi,t}$. The plot is taken over the spherical contour $r=0.75\,\text{au}$ and has been averaged in azimuth and over the interval $t_{\text{in}}\in[200,700]$. There are strong localised accretion flows (grey shaded regions) in the disc surface layers, which are associated with radial angular momentum transport of a magnetic origin, quantified in Fig. 8.
@@
To understand the vertical morphology in the active zone, we examine the magnetic field structures and their associated stresses over the spherical contour, $r=0.75$ au. The $(\theta,t)$ spacetime diagram presented in Fig. 6 exhibits the operation of the sustained stratified MRI dynamo close to the disc midplane throughout the entire simulation, alongside the associated butterfly pattern in $\langle B_{\phi}\rangle_{\phi}$ which corresponds to periodic azimuthal field cycles (e.g. Brandenburg et al., 1995).
@@
However, close to the disc surface layers there is a different phenomenon, responsible for driving surface localised accretion. The strong build-up of radial and toroidal field dramatically alters the vertical structure of magnetic fields in the active region. First, the corresponding plasma-$\beta$ parameter (24) along this contour, varies from $\sim 10^{4}$ at the disc midplane, down to $\sim 10$ at $z=\pm 3H$. Whilst, this is not sufficient to suppress the operation of the MRI through magnetic tension, it indicates the presence of two different magnetic field strength regimes. Meanwhile, local magnetic field correlations drive a sustained $\alpha_{\mathcal{M}}\sim 10^{-1}$ in the disc surface layers, in sharp contrast to the weak $\alpha_{\mathcal{M}}\sim 10^{-3}$ at the disc midplane, as shown in Fig. 6.
@@
In Fig. 7, we present the resulting layered accretion profile, characterised by $\langle\rho u_{r}\rangle_{\phi,t}$ along the same spherical contour. The averaged accretion rate is an order of magnitude stronger (grey shaded regions) in the surface layers compared to the disc midplane. Whilst this accretion flow is not supersonic, it reaches approximately 75 per cent of the local disc sound speed, $c_{s}=\varepsilon R^{-1/2}$, at the altitude of the maximal accretion rate. This bears resemblance to the elevated accretion structures observed in vertical-net flux global MRI active simulations performed by Zhu & Stone (2018) and Jacquemin-Ide et al. (2021).
@@
Finally, we decompose the vertical profiles of these stresses further in Fig. 8 to take account of the turbulent and coherent components, defined in equation (23). Even though vertical stresses contribute $\sim\!R/H$ more to accretion than vertically integrated radial stresses, under the thin-disc approximation, the dominant driver of outward angular momentum transport is the turbulent $r\phi$ component of the Maxwell stress, $\mathcal{M}^{\text{turb}}_{r\phi}$. Therefore, despite the large-scale, laminar looking magnetic field structures (see Fig. 3), the vertical accretion structure in the MRI active zone is being driven by significant vertical inhomogeneity in the radial torque imposed by $\mathcal{M}^{\text{turb}}_{r\phi}$.
@@
Figure 8: Decomposition of the total Maxwell stress components, defined in equation (23), along the spherical contour $r=0.75\,\text{au}$. The stresses are normalised by the local pressure to enable quantitative comparisons across altitudes, and have been averaged in azimuth and over the interval $t_{\text{in}}\in[200,700]$. The grey shaded regions indicate the localised surface accretion flows at $z\approx\pm 3H$, shown in Fig. 7.
@@
To test the sensitivity of surface layer accretion in the active region to the numerical boundary proximity and the dead–active zone interface, we conducted an additional simulation, ZNF-AZ (see Table 3.3), where the non-ideal MHD diffusivity profiles are removed, rendering the entire disc fully MRI turbulent. In this model, there is also the accumulation of a strong, tightly wound toroidal magnetic field just beneath the disc’s surface layers, as shown in the $(\theta,t)$ spacetime diagrams in Fig. 9. This should be compared with the corresponding contour from the ZNF-FID model in Fig. 6. Moreover, the configuration extends across the full radial extent of the disc, including regions well beyond the inner boundary. This suggests that the behaviour is not solely driven by boundary conditions or the imposed dead–active zone interface, but is intrinsic to the ZNF MRI and thermodynamic setup used. Further work, in a more controlled environment, is needed to determine the exact mechanism that forms this magnetic field morphology. For now, we hypothesise the following general argument.
@@
Figure 9: Spacetime $(\theta,t)$ diagrams for the fully MRI active disc simulation, ZNF-AZ, along the contour $r=0.75$ au for the azimuthally averaged toroidal field $\langle B_{\phi}\rangle_{\phi}$ (top) and Maxwell $\alpha$ component (bottom). The disc–corona temperature transition is at $z=\pm 4H$ (dashed black lines) and these results should be compared with the similar plots for ZNF-FID presented in Fig. 6.
4.3.2 Surface layer accretion mechanism
@@
The sharp transition to a hot corona, defined in equation (7), alters the buoyancy forces that push the dynamo generated magnetic flux away from the disc midplane. In the simplified one-dimensional flux tube model of Torkelsson (1993), a sharp gradient change in the vertical pressure profile can stall buoyantly rising azimuthal flux tubes of fixed length at the disc–corona transition, resulting in the build up of these magnetic flux tubes just below the disc surface layers.
@@
In our simulations, the large-scale poloidal magnetic field loops fill the active zone, but do not expand vertically into the corona. Instead, the generated large-scale radial field is halted at the disc–corona transition, as evidenced by the radial flux bunching at $|z|\approx 0.2$ in Fig. 5. We attribute this behaviour to the modified buoyancy profile at the transition, but note that it represents a more complex configuration than that considered in the simplified model of Torkelsson (1993) discussed above. The $\Omega$-effect acts on the buoyantly trapped radial field, accessing the disc’s shear reservoir to generate strong toroidal fields at both surface layers, as supported by the observed anti-correlation between the signs of $B_{r}$ and $B_{\phi}$ in Fig. 5. Once formed, the large-scale poloidal loops (which are always clockwise in our simulations) stabilise a particular polarity, effectively locking in the sign of $B_{\phi}$ produced by the $\Omega$-effect for the remainder of the evolution.
@@
Additional evidence for the influence of a hot corona on the structure of MRI active discs is found in previous studies of both ZNF and vertical-net flux simulations. For example, the shearing box simulations by Bambic et al. (2024), with a self-consistently computed two-temperature disc–corona structure demonstrate similar features (see bottom panels of their fig. 1). Comparable behaviour is also seen in global simulations of vertical-net flux presented in the second paper of this series and in ZNF simulations by J. Jacquemin-Ide (private communication). In contrast, the global ZNF MRI active simulations by Parkin & Bicknell (2013) show no such evidence, despite implementing a disc–corona temperature transition (see their fig. 7). However, their method imposes the transition using a different, possibly less sharp, technique and employs a geometrically thinner disc ($\varepsilon=0.05$), a regime that has previously raised concerns about the generation of large-scale poloidal fields from the MRI dynamo (Jacquemin-Ide et al., 2024).
4.4 Midplane accretion structure in the disc

### v3

@@ 4.2.1 Emergence and saturation of the MRI
Figure 4: Temporal evolution of the volume-averaged total $\alpha$ (black) and Maxwell $\alpha_{\mathcal{M}}$ (red) in the MRI-active zone, which are larger in the higher-resolution simulation ZNF-HRES (dotted) compared to ZNF-FID (solid). The volume average is computed over the spherical wedge: $r\in[0.5,1]\,\text{au}$, $\theta\in[z=-4H,z=4H]$ and $\phi\in[0,2\pi)$.
We begin by characterising the saturated MRI dynamo in the ideal-MHD region of the disc. As shown in Fig. 4, the volume-averaged $\alpha$ in the ideal-MHD zone exhibits rapid exponential growth, reaching saturation within approximately $3\,t_{\mathrm{DZI}}$, and remains sustained for the duration of the simulation. The time-averaged value is $\alpha\!\sim\!8\times 10^{-3}$, with the Maxwell component contributing $\alpha_{\mathcal{M}}\!\sim\!5\times 10^{-3}$. This behaviour confirms the presence of a persistent MRI-driven dynamo operating via a fully developed turbulent state. For comparison, the global ZNF dead–active zone interface simulations with radiative transfer by flock_3d_2017 report a higher volume-averaged stress of $\alpha\!\sim\!3\times 10^{-2}$ in the MRI-active zone of their fiducial model. However, their azimuthally restricted domain employed a higher effective resolution, equivalent to 2048 cells in azimuth.
To assess the sensitivity of this evolution to the numerical grid, we conduct a resolution test using ZNF-AZ. The disc structure, evolution, and large-scale magnetic features resemble those in ZNF-FID, but differences arise due to changes in the saturated strength of the ZNF MRI. As shown in Fig. 4, the total and Maxwell components of the saturated, volume-averaged $\alpha$ in the active zone, are about $60$ per cent higher in ZNF-HRES (dotted lines), enhancing angular momentum transport and accelerating viscous evolution by a similar factor, under thin-disc assumptions. This underscores the difficulty of extracting precise accretion rates from global ZNF MRI simulations at these resolutions. Whether higher-resolution global simulations reproduce the non-converging behaviour in stratified isothermal shearing boxes (see table 1 in ryan_resolution_2017), where $\alpha$ falls with increasing resolution, remains unclear.
@@ 4.2.2 Large-scale coherent poloidal magnetic field loops
In the MRI-active region, large-scale poloidal field loops are generated self-consistently from an initial toroidal field, which is defined in equation (12), and grow to fill the extent of the active disc region. Fig. 5 shows the magnetic field configuration in the inner disc, averaged in azimuth and over the time interval $t_{\mathrm{in}}\in[200,700]$, to enhance the visibility of the coherent field structures against the background turbulent state. The poloidal field configuration, $\langle\mathbf{B_{p}}\rangle_{\phi,t}$, is visualised using field lines (black) and a LIC overlay (greyscale). A set of concentric, clockwise, large-scale poloidal field loops is clearly visible. Once formed, these structures persist throughout the extent of the simulation (see Section 6), and their comparable evolution in the azimuthally restricted, long-runtime simulation, ZNF-LONG, indicates that a full azimuthal domain is not required to capture this process.
Previous work by jacquemin-ide_magnetorotational_2024 demonstrated the self-consistent generation of large-scale poloidal magnetic field loops in global MRI-active simulations of black hole accretion discs. In contrast to our initial conditions (see Section 3.3.1), their study began with a strong, uniform axisymmetric toroidal magnetic field with a plasma-beta of $\beta\!=\!5$ and a scale height of $H/r\!\sim\!0.2$ in the bulk disc, corresponding to a relatively thick accretion torus structure. However, they raised concerns regarding the efficiency of dynamo-driven field generation in thinner discs, which are less effective at advecting magnetic flux (lubow_magnetic_1994). Therefore, this work provides evidence for the formation and persistence of similar large-scale poloidal field loops in a thinner disc with $H/R=0.1$, a regime more representative of protoplanetary discs. This lends tentative support to the notion that thin MRI-active discs can generate and sustain large-scale magnetic fields self-consistently, without requiring an initial vertical-net flux.
Figure 5: Magnetic field configuration for ZNF-FID in the active zone, averaged in azimuth and over the interval $t_{\text{in}}\in[200,700]$ to enhance the visibility of coherent field structures. The poloidal magnetic field, $\langle\mathbf{B_{p}}\rangle_{\phi,t}$, shown using black field lines and a LIC overlay in greyscale, forms coherent clockwise loops spanning the extent of the MRI-active region. The sign of the normalised toroidal field, $R\langle B_{\phi}\rangle_{\phi,t}$, (background colour) near the disc surface layers (dashed white lines), is negatively correlated with the sign of the radial magnetic field.
@@ 4.3 Vertical structure of the active zone
4.3.1 Surface-layer accretion
@@
In the MRI-active, ideal-MHD region of the disc a vertically layered accretion structure develops, and is sustained, in all the simulations, including the pure active model ZNF-AZ. To our knowledge, this feature has not previously been reported in the ZNF configuration. The majority of the accretion is localised just below the disc surface layers at $z\approx\pm 3H$, in a near symmetric manner across the disc midplane. To quantify the connection between accretion and magnetic stresses, we perform an extensive temporal average over the interval $t_{\mathrm{in}}\in[200,700]$ in this section.
@@
Figure 6: Spacetime $(\theta,t)$ diagrams for ZNF-FID, along the contour $r=0.75$ au, showing the azimuthally averaged toroidal field, $\langle B_{\phi}\rangle_{\phi}$ (top), and Maxwell $\alpha$ component (21) (bottom). The disc–corona transition is at $z=\pm 4H$ (dashed black lines) and the surface layers above the MRI-active disc have a tightly wound magnetic field configuration that drives strong local radial angular momentum transport.
@@
Figure 7: Vertical structure of the accretion flow in the MRI-active region for ZNF-FID (blue) and ZNF-AZ (red), quantified by $\langle\rho u_{r}\rangle_{\phi,t}$. The plot is taken over the spherical contour $r=0.75\,\text{au}$ and has been averaged in azimuth and over the interval $t_{\text{in}}\in[200,700]$. There are strong localised accretion flows (grey-shaded regions) in the disc surface layers, which are associated with radial angular momentum transport of a magnetic origin, quantified in Fig. 8.
@@
To understand the vertical morphology in the active zone, we examine the magnetic field structures and their associated stresses over the spherical contour $r=0.75$ au. The $(\theta,t)$ spacetime diagram presented in Fig. 6 exhibits the operation of the stratified MRI dynamo close to the midplane throughout the entire simulation, showing the well-known butterfly pattern in $\langle B_{\phi}\rangle_{\phi}$, which corresponds to periodic azimuthal-field cycles (e.g. brandenburg_dynamo-generated_1995).
@@
However, close to the disc surface layers there is a different phenomenon, responsible for driving surface-layer localised accretion. The strong build-up of radial and toroidal field dramatically alters the vertical structure of magnetic fields in the active region. First, the corresponding plasma-$\beta$ parameter, defined in (23), along this contour, varies from $\sim\!10^{4}$ at the midplane, down to $\sim\!10$ at $z=\pm 3H$. Whilst this is not sufficient to suppress the MRI through magnetic tension, it indicates the presence of two different magnetic field strength regimes. Meanwhile, local magnetic field correlations drive a sustained $\alpha_{\mathcal{M}}\!\sim\!10^{-1}$ in the disc surface layers, in sharp contrast to the weak $\alpha_{\mathcal{M}}\!\sim\!10^{-3}$ at the midplane, as shown in Fig. 6.
@@
In Fig. 7, we present the resulting vertically layered accretion profile, characterised by $\langle\rho u_{r}\rangle_{\phi,t}$, along the same spherical contour for both ZNF-FID and ZNF-AZ. The averaged accretion rate is an order of magnitude stronger (grey-shaded regions) in the surface layers compared to the midplane in ZNF-FID. Whilst this accretion flow is not supersonic, it reaches approximately 75 per cent of the local disc sound speed, $c_{\text{s}}=\varepsilon R^{-1/2}$, at the altitude of the maximal accretion rate. This bears resemblance to the elevated accretion structures observed in vertical-net flux global MRI-active simulations performed by zhu_global_2018 and jacquemin-ide_magnetic_2021.
@@
Finally, we decompose the vertical profiles of these stresses further in Fig. 8 to take account of the turbulent and coherent components, defined in equation (22). Even though vertical stresses contribute $\sim\!R/H$ more to accretion than vertically integrated radial stresses, under the thin-disc approximation, the dominant driver of outward angular momentum transport is the turbulent $r\phi$ component of the Maxwell stress, $\mathcal{M}^{\text{turb}}_{r\phi}$. Therefore, despite the large-scale, laminar looking magnetic field structures (see Fig. 3), the vertical accretion structure in the MRI-active zone is being driven by significant vertical inhomogeneity in the radial torque imposed by $\mathcal{M}^{\text{turb}}_{r\phi}$.
@@
Figure 8: Decomposition of the total Maxwell stress components, defined in equation (22), along the spherical contour $r=0.75\,\text{au}$. The stresses are normalised by the local pressure to enable quantitative comparisons across altitudes, and have been averaged in azimuth and over the interval $t_{\text{in}}\in[200,700]$. The grey-shaded regions indicate the localised surface accretion flows at $z\approx\pm 3H$, shown in Fig. 7.
@@
To test the sensitivity of surface-layer accretion in the active region to the numerical boundary proximity and the dead–active zone interface, we conducted an additional simulation, ZNF-AZ (see Table 3.3), where the non-ideal MHD diffusivity profiles are removed, rendering the entire disc fully MRI turbulent. In this model, there is also the accumulation of a strong, tightly wound toroidal magnetic field just beneath the disc’s surface layers, as shown in the $(\theta,t)$ spacetime diagrams in Fig. 9. This should be compared with the corresponding contour from the ZNF-FID model in Fig. 6. Moreover, the configuration extends across the full radial extent of the disc, including regions well beyond the inner boundary. This suggests that the behaviour is not solely driven by boundary conditions or the imposed dead–active zone interface, but is intrinsic to the ZNF MRI and thermodynamic setup used. Further work, in a more controlled environment, is needed to determine the exact mechanism that forms this magnetic field morphology. For now, we hypothesise the following general argument.
@@
Figure 9: Spacetime $(\theta,t)$ diagrams for the fully MRI-active disc simulation, ZNF-AZ, along the contour $r=0.75$ au for the azimuthally averaged toroidal field, $\langle B_{\phi}\rangle_{\phi}$ (top), and Maxwell $\alpha$ component (bottom). The disc–corona temperature transition is at $z=\pm 4H$ (dashed black lines) and these results should be compared with equivalent plots for ZNF-FID presented in Fig. 6.
4.3.2 Surface-layer accretion mechanism
@@
The sharp transition to a hot corona, defined in equation (7), alters the buoyancy forces that push the dynamo-generated magnetic flux away from the disc midplane. In the simplified one-dimensional flux-tube model of torkelsson_magnetic_1993, a sharp gradient change in the vertical pressure profile can stall buoyantly rising azimuthal flux tubes of fixed length at the disc–corona transition, resulting in the build-up of these magnetic flux tubes just below the disc surface layers.
@@
In our simulations, the large-scale poloidal magnetic field loops fill the active zone, but do not expand vertically into the corona. Instead, the generated large-scale radial field is halted at the disc–corona transition, as evidenced by the bunching of radial flux at $|z|\approx 0.2\,\text{au}$ in Fig. 5. We attribute this behaviour to the modified buoyancy profile at the transition, but note that it represents a more complex configuration than the simplified model of torkelsson_magnetic_1993 discussed above. The $\Omega$-effect acts on the buoyantly trapped radial field, accessing the disc’s shear reservoir to generate strong toroidal fields at both surface layers, as supported by the observed anti-correlation between the signs of $B_{r}$ and $B_{\phi}$ in Fig. 5. Once formed, the large-scale poloidal loops (which are always clockwise in our simulations) stabilise a particular polarity, effectively locking in the sign of $B_{\phi}$ produced by the $\Omega$-effect for the remainder of the evolution.
@@
Additional evidence for the influence of a hot corona on the structure of MRI-active discs is found in previous studies of both ZNF and vertical-net flux simulations. For example, the shearing-box simulations by bambic_local_2024, with a self-consistently computed two-temperature disc–corona structure demonstrate similar features (see bottom panels of their fig. 1). Comparable behaviour is also seen in global simulations of vertical-net flux presented in the second paper of this series and in ZNF simulations by J. Jacquemin-Ide (private communication). In contrast, the global ZNF MRI-active simulations by parkin_global_2013 show no such evidence, despite implementing a disc–corona temperature transition (see their fig. 7). However, their method imposes the transition using a different, possibly less sharp, technique and employs a geometrically thinner disc ($\varepsilon=0.05$), a regime that has previously raised concerns about the generation of large-scale poloidal fields from the MRI dynamo (jacquemin-ide_magnetorotational_2024).
4.4 Midplane accretion structure